Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2015-02-12

Observers : Sylvain Baumont, Paula Zelaya
TAT : Joe Anderson, Nancy Elias-Rosa

We did mainly follow-up observations with EFOSC, and classified two SN-Ia.

Clear sky, low wind, below 50% humidity.
Photometric night.

(times are in UT)

CALIBRATIONS :
 - Evening : EFOSC_bias, EFOSC_FlatSpec, EFOSC_SpecArc
 - 22:00 : Dome OPEN
 - 23:56 : Attempt to get EFOSC_SkyFlat
    - 00:05  "NTT tracking control" issue => failed to complete SkyFlat sequence.
    - 00:16  It's official, "WE HAVE A PROBLEM !"
 - 09:xy : EFOSC_SkyFlat

SCIENCE OBSERVATIONS :
--- EFOSC ---

 - 00:48  Attempt to get STD GD71 : EFOSC "severe error"
               The rotator encoder failed due to heat. Shut down for 30'. Reboot.
    - 02:14  EFOSC is back ! (2h lost, but the seeing was >1" :P )
                  No image analysis is needed, from Acq image.

 - 02:15  LSQ14gqk FOLLOW-UP : Gr13/3x1500s, slit 1", airmass 1.4-1.7, seeing 1.0-1.4"

 - 03:45  Photometric Standard Field PG1047 : UBVRigrz, airmass 1.37, seeing 0.95"

 - 04:01  LSQ14efd FOLLOW-UP : B/3x300s, VRi/3x200s, airmass 1.9-2.3, seeing 0.8"

 - 04:58  LSQ15im CLASSIFICATION : Gr13/900s, slit 1", airmass 1.4, seeing 0.8"
               => SN-Ia +1week, z ~ 0.075

 - 05:14  Photometric Standard Field RU152 : UBVRigrz, airmass 1.43, seeing 0.8"

 - 05:26  MLS150210-102122+110418 CLASIFICATION : Gr13/1500s, slit 1", airmass 1.31, seeing 0.8"
               => SN-Ia ~max, z ~ 0.15

 - 06:01  STD L745a : Gr11, Gr13, Gr16

 - 06:12  PSNJ09204691-0803340 FOLLOW-UP : B/3x300, VRi/3x200s, airmass 1.2, seeing 0.9"
              (Manual offset to avoid bad column at SN location)
 
 - 07:20  ASASSN-15ba FOLLOW-UP : Gr11/2x1500s, Gr16/2x1500s, slit 1", airmass 1.4, seeing 0.8"

 - 09:20  Photometric Standard Field PG1657 : UBVRigrz, airmass 1.74, seeing 0.7"

CONCLUSIONS :

LSQ15im  :  SN-Ia +1week, z ~ 0.075
MLS150210-102122+110418  :  SN-Ia around max, z ~ 0.15
 

Comments